# On the Need to Rethink Trust in AI Assistants for Software Development: A Critical Review

### v2

@@ III-B2 Human-Computer Interaction (HCI)
Trust has been studied in HCI since the late 1980s [103]. Our review of HCI literature includes articles published between 1996 and 2024. Among the 24 unique HCI research papers we included, eight articles embedded their trust definitions in [95]’s model. Three articles explicitly cite the definition, one paraphrases it, and four reference components such as ability, integrity, or benevolence without adopting the complete model. Furthermore, nine papers refer to [80]’s trust model, and another five papers incorporate elements from both models. Among the twelve papers building upon existing trust models, six extended the models either by developing new models or by adopting the definitions and further exploring specific aspects of trust. The remaining six papers refer to the models without modification. In contrast, eleven papers do not explicitly refer to existing trust models.
@@
Several HCI articles conceptualize trust as a context-dependent phenomenon shaped by system behavior, user interaction, and situational factors. This perspective is especially prevalent in domains such as AI agents, automated decision support systems, and human-robot interaction. For example, [139] [139] propose the human-centered trustworthy framework, which adapts [52]’s human–computer trust model [52] and integrates trustworthiness components from [95] [95], along with constructs from [98] [98] and [80] [80]. In this framework, trust is approached as a multilevel concept that includes user predispositions, interface features such as transparency and control, and institutional factors such as data governance and accountability. This layered view underscores the relational and evolving nature of trust in human–AI interaction.
@@
[114] [114] developed and validated a human–robot interaction trust scale by adapting [52]’s model [52]. Trust is conceptualized following [95] [95]’s definition as the willingness to be vulnerable to another party, and is further informed by [80]’s view of trust development over time [80]. Their adapted model incorporates dimensions such as reciprocity, competence, benevolence, predictability, honesty, trust predisposition, and willingness to be vulnerable, tailored to human–robot interaction.
@@
[109] [109] conceptualize trust as a multidimensional and dynamic construct that evolves through interaction. They draw on [95] [95] to describe trustworthiness in terms of ability, integrity, and benevolence, and incorporate [80]’s view of trust in automation [80] as a calibrated attitude shaped by system performance and user experience. Their conceptualization also reflects [90]’s adaptation [90] of [96]’s definition [96], which emphasizes user confidence and willingness to act on the recommendations of intelligent systems. Building on this foundation, [109] [109] propose an adaptive explanation architecture that adjusts system transparency and feedback to support appropriate levels of trust in automated systems.
@@
[117] [117] conceptualize trust as a psychological state characterized by a positive expectation that one’s vulnerability will not be exploited. Like [109] [109], they build on [95] [95] and [96] [96], but extend the discussion by emphasizing the social grounding of trust. Their framework is embedded in a trustor–trustee model adapted from the trust game [9], where trust is shaped by contextual cues, reputation, and internalized norms. They argue that trust in mediated interactions is not only influenced by individual traits but also by perceived trustworthiness of systems shaped through interface design and social context.
@@
In contrast, several papers reference foundational trust models without adapting or extending them. [15] [15] cite [95] [95], [96] [96] and [90] [90] to examine how untrustworthy AI systems can still elicit user trust, but do not propose a new framework or modify the referenced models. Similarly, [165] [165] apply [80] [80] definition to investigate trust in AI-based medical decision-making, without extending it conceptually. [142] [142] adopt [80]’s trust model to frame user trust in AI-based video interviews. Although the model is applied to a specific domain, its theoretical structure remains unchanged. [78] [78] primarily build on [90] [90] trust model, while only briefly mentioning [95] and [80]. These models are not embedded in their conceptual framework and rather serve as peripheral citations. Similarly, [28] [28] adopt [95]’s and [80]’s definitions of trust to conceptualize user trust in human–computer interaction, but use the models primarily to validate the Chinese version of the human–computer trust scale in a new cultural context.
@@ III-B3 Software Engineering (SE)
Trust is a critical element in SE that influences the adoption of software tools, processes, and collaborative practices. However, our review of 13 articles published between 2002 and 2024 reveals that SE research typically does not engage with established foundational trust models. In fact, only one paper [63] cited [95]’s trust definition and based their research on a related trust model. Furthermore, one paper [155] implicitly referred to established trust models by mentioning risk, that is, the aspect of vulnerability (see Section II-A). The other papers either did not explicitly define trust or used custom definitions, diverging from disciplines such as IS and HCI, where established trust models are more commonly applied.
An example of SE’s customized approach to trust is the PICSE framework proposed by [66] [66]. It describes a structured approach to understanding trust in software tools, identifying five key factors: personal, interaction, control, system, and expectation. The authors refer to [42]’s credibility framework [42] from HCI and [118]’s trust literature review [118] from IS. However, these trust models did not contribute to PICSE but are discussed after introducing the framework.
[63] adapted [130]’s situational trust model [130], which extends [95]’s model, to analyze trust dynamics in global software engineering [63]. In line with HCI and IS research, their model underlines the dynamic nature of trust, distinguishing initial trust building, where trust forms based on past interactions, from trust evolution, where trust fluctuates during the project based on expectations and behaviors.
[155] applied evolutionary game theory to examine trust and cooperation in distributed teams. Their study does not explicitly model trust development or draw on established trust theories [155]. Similarly, [29] propose dynamic assurance cases for engineering trustworthy self-adaptive software, emphasizing system reliability and correctness rather than interpersonal trust [29]. [3] take a different approach, presenting a taxonomy of decision-making challenges in trustworthy AI [3]. Although these studies address trust-related concerns, they do so in domain-specific ways, emphasizing technical, ethical, or structural factors rather than explicitly modeling trust development based on established theories.
[23] address trust in single data source software engineering case studies [23]. [20] investigate trust perceptions in code review through an eye-tracking study, demonstrating how the perceived origin of code (human versus machine-generated) influences review behaviors and trust judgments [20]. However, both papers do not explicitly reference or build upon established trust models.
@@ III-B4 Literature Reviews and Special Issues
We found five literature reviews of trust that were not limited to one of our target disciplines SE, HCI, or IS. [71]’s review centers around requirements for trustworthy AI [71]. They refer to trust definitions across disciplines, including psychology, sociology, and economy, mentioning integrity and reliability as an agreement across these disciplines. They define trustworthy AI as a framework to ensure that a system is worthy of being trusted based on the evidence concerning its stated requirements [71]. [163] examined 142 studies published between 2015 and 2022, identifying components, influencing factors, and outcomes of users’ trust in AI [163]. [81]’s review outlines aspects of AI trustworthiness such as robustness, explainability, transparency, and fairness, organized along the lifecycle of AI systems [81]. However, they do not discuss established trust definitions or models. [51] reviewed 47 studies on trust in technology published in IS and HCI venues and found 17 different theories that were integrated from disciplines such as psychology and economics [51]. They conclude that the intricacies of how trust is formed and maintained in online environments still necessitates further investigation and that the development of standardized and empirically validated instruments for trust measurement is crucial. Finally, [88]’s review focuses on antecedents of trust in AI [88]. They identify five AI-specific trust challenges: transparency and explainability, accuracy and reliability, automation versus augmentation, anthropomorphism and embodiment, and mass data extraction [88]. An interesting observation regarding AI assistants for software development is that over-anthropomorphism may lead to overestimation of the AI’s capabilities [88], increasing risk [34], decreasing trust [11], and potentially leading to manipulation [123]. This is certainly an underexplored perspective on AI assistants for software development, especially when focusing on human factors of software security.
@@
Three literature reviews focused on trust research in HCI. [65]’s review discussed the effects of emotions on trust in the context of technology use [65]. Their review indicates that positive emotions lead to higher trust. [8] reviewed 23 empirical studies that focus on user trust in AI-enabled systems [8]. They describe trust definitions, factors that influence user trust, and measurement methods, reporting that seven studies explicitly defined trust, six of which referring to [95] [95] or [80] [80]. Eight studies conceptualized trust, while nine studies neither defined nor conceptualized trust. They found surveys, interviews, and focus groups to be the most commonly used methods to assess user trust. [8] conclude that it is important to select a definition of user trust that aligns with the specific study context. [102] reviewed 65 articles to assess the various definitions of appropriate trust in human-AI interaction, providing an overview of concepts and definitions and identifying similarities and differences between them [102]. They conclude that four common methods for building appropriate trust are improving system transparency, considering user cognition and perception, using guidelines or models to achieve calibrated trust in AI, and considering the entire continuum of trust, including over-, under-, mis-, and dis-trust.
@@
One literature review had a clear IS focus. [49] synthesized two decades of empirical research on the determinants of human trust in AI across disciplines [49]. The authors motivate how trust in AI differs from other technologies and identify an AI’s capabilities as an important antecedent of trust development. In addition, they propose a framework that identifies factors that influence users’ cognitive and emotional trust. Regarding the development of cognitive trust, they highlight the importance of an AI’s tangibility, transparency, reliability, and immediacy. For emotional trust, on the other hand, and AI’s anthropomorphism is a major factor. [49] also critically reflect on past studies by pointing to the diversity of trust measure used and the lack of longer-term studies in higher stakes environments [49].
@@
For IS, we further identified three special issues on trust. One of them was published in 2010 in their flagship journal MIS Quarterly, focusing on novel perspectives on trust in IS. The fact that this special issue was published 15 years ago is another indication of the maturity of trust research in that discipline. The other IS-specific special issues focused on trust in online environments [18] and trust in AI [89].
@@
For SE, we included two literature reviews. [86] conducted a tertiary study reviewing 141 secondary studies on trustworthy AI and software, describing trustworthiness as a highly abstract concept comprising related quality attributes [86]. The review focuses on these quality attributes and compares them between trustworthy AI and other trustworthy software. [86] do not focus on the conceptualization of trust and trustworthiness. [59] reviewed 112 articles on trust in software ecosystems, identifying different definitions of concepts such as software trust, system trust, or software service trust. However, their focus was on compiling trust definitions, not unifying or aligning them [59].
@@ III-C Critical Review of Trust Conceptualizations in Software Engineering Research
Of the 58 articles that we analyzed as part of our critical review, only one article referred to an established trust model, and only three articles provided trust definitions. This underlines the problematic situation suggested by our literature across disciplines (see Section III-B3). There is a significant gap in the maturity of trust research in SE compared to IS and HCI. Trust as a concept is often discussed in a context-specific manner without embedding concepts and definitions in established trust theories or models. This lack of embedding in established trust theories from other disciplines hinders interdisciplinary comparisons and research towards a deeper understanding of what constitutes trust in SE-specific settings.
@@
[108] are the only authors who explicitly reference an established trust model, incorporating [80]’s model of human-automation trust [108]. [4] define trust as the unyielding belief that the person is truthful and reliable and in the context of their research as the willingness of the community to rely on the contributor, a prerequisite for considering her code change [4]. This conceptualization frames trust as a social and reputation-based construct, essential in collaborative environments such as open-source software development, where trust is built through ongoing contributions. Although this framing is reasonable, an embedding in existing trust models would allow researchers to identify the particularities of trust in open-source software compared to other settings.
@@
[158] [158] implicitly define trust by referencing dispositional trust as proposed by [58] [58], which describes an individual’s inherent tendency to trust automation, independent of context or system performance. This positions trust as a psychological trait that varies between individuals. However, it is important to note that trust was only a peripheral topic, not a central focus of their study. [77] discuss trust in the context of large language models (LLMs) for code generation, suggesting that users assess trustworthiness based on the model’s explainability and how well its predictions align with their expectations [77]. However, they neither explicitly define trust nor consider the broader explainability discussion (see Section II). Instead, they cite previous work by [22], who explored how users determine whether to trust a machine learning model based on its ability to convey behavior in a way that aligns with their expectations [22]. Neither study provides a formal definition of trust; rather, both frame it in terms of user perceptions and interpretability. Given the amount of related work on trustworthiness and explainability of AI-based systems in philosophy (see Section II-C) and IS (see Section III-B1), it is surprising that this article is again not broadly embedded in established trust conceptualizations.

### v3

@@ III-B2 Human-Computer Interaction (HCI)
Trust has been studied in HCI since the late 1980s [DBLP:journals/ijmms/Muir87]. Our review of HCI literature includes articles published between 1996 and 2024. Among the 24 unique HCI research papers we included, eight articles embedded their trust definitions in [Mayer1995]’s model. Three articles explicitly cite the definition, one paraphrases it, and four reference components such as ability, integrity, or benevolence without adopting the complete model. Furthermore, nine papers refer to [Lee2004Trust]’s trust model, and another five papers incorporate elements from both models. Among the twelve papers building upon existing trust models, six extended the models either by developing new models or by adopting the definitions and further exploring specific aspects of trust. The remaining six papers refer to the models without modification. In contrast, eleven papers do not explicitly refer to existing trust models.
@@
Several HCI articles conceptualize trust as a context-dependent phenomenon shaped by system behavior, user interaction, and situational factors. This perspective is especially prevalent in domains such as AI agents, automated decision support systems, and human-robot interaction. For example, [DBLP:journals/computer/SousaLCM24] [DBLP:journals/computer/SousaLCM24] propose the human-centered trustworthy framework, which adapts [DBLP:journals/behaviourIT/Gulati0L19]’s human–computer trust model [DBLP:journals/behaviourIT/Gulati0L19] and integrates trustworthiness components from [Mayer1995] [Mayer1995], along with constructs from [DBLP:journals/tmis/McKnightCTC11] [DBLP:journals/tmis/McKnightCTC11] and [Lee2004Trust] [Lee2004Trust]. In this framework, trust is approached as a multilevel concept that includes user predispositions, interface features such as transparency and control, and institutional factors such as data governance and accountability. This layered view underscores the relational and evolving nature of trust in human–AI interaction.
@@
[pinto2022trust] [pinto2022trust] developed and validated a human–robot interaction trust scale by adapting [DBLP:journals/behaviourIT/Gulati0L19]’s model [DBLP:journals/behaviourIT/Gulati0L19]. Trust is conceptualized following [Mayer1995] [Mayer1995]’s definition as the willingness to be vulnerable to another party, and is further informed by [Lee2004Trust]’s view of trust development over time [Lee2004Trust]. Their adapted model incorporates dimensions such as reciprocity, competence, benevolence, predictability, honesty, trust predisposition, and willingness to be vulnerable, tailored to human–robot interaction.
@@
[DBLP:conf/compsac/NothdurftBLM12] [DBLP:conf/compsac/NothdurftBLM12] conceptualize trust as a multidimensional and dynamic construct that evolves through interaction. They draw on [Mayer1995] [Mayer1995] to describe trustworthiness in terms of ability, integrity, and benevolence, and incorporate [Lee2004Trust]’s view of trust in automation [Lee2004Trust] as a calibrated attitude shaped by system performance and user experience. Their conceptualization also reflects [madsen2000measuring]’s adaptation [madsen2000measuring] of [mcallister1995affect]’s definition [mcallister1995affect], which emphasizes user confidence and willingness to act on the recommendations of intelligent systems. Building on this foundation, [DBLP:conf/compsac/NothdurftBLM12] [DBLP:conf/compsac/NothdurftBLM12] propose an adaptive explanation architecture that adjusts system transparency and feedback to support appropriate levels of trust in automated systems.
@@
[DBLP:journals/ijmms/RiegelsbergerSM05] [DBLP:journals/ijmms/RiegelsbergerSM05] conceptualize trust as a psychological state characterized by a positive expectation that one’s vulnerability will not be exploited. Like [DBLP:conf/compsac/NothdurftBLM12] [DBLP:conf/compsac/NothdurftBLM12], they build on [Mayer1995] [Mayer1995] and [mcallister1995affect] [mcallister1995affect], but extend the discussion by emphasizing the social grounding of trust. Their framework is embedded in a trustor–trustee model adapted from the trust game [2001-16661-005], where trust is shaped by contextual cues, reputation, and internalized norms. They argue that trust in mediated interactions is not only influenced by individual traits but also by perceived trustworthiness of systems shaped through interface design and social context.
@@
In contrast, several papers reference foundational trust models without adapting or extending them. [DBLP:journals/pacmhci/BanovicYRL23] [DBLP:journals/pacmhci/BanovicYRL23] cite [Mayer1995] [Mayer1995], [mcallister1995affect] [mcallister1995affect] and [madsen2000measuring] [madsen2000measuring] to examine how untrustworthy AI systems can still elicit user trust, but do not propose a new framework or modify the referenced models. Similarly, [DBLP:journals/ijhci/YokoiEFN21] [DBLP:journals/ijhci/YokoiEFN21] apply [Lee2004Trust] [Lee2004Trust] definition to investigate trust in AI-based medical decision-making, without extending it conceptually. [DBLP:journals/chb/SuenH23] [DBLP:journals/chb/SuenH23] adopt [Lee2004Trust]’s trust model to frame user trust in AI-based video interviews. Although the model is applied to a specific domain, its theoretical structure remains unchanged. [DBLP:conf/um/KrausWM20] [DBLP:conf/um/KrausWM20] primarily build on [madsen2000measuring] [madsen2000measuring] trust model, while only briefly mentioning [Mayer1995] and [Lee2004Trust]. These models are not embedded in their conceptual framework and rather serve as peripheral citations. Similarly, [Cai02012025] [Cai02012025] adopt [Mayer1995]’s and [Lee2004Trust]’s definitions of trust to conceptualize user trust in human–computer interaction, but use the models primarily to validate the Chinese version of the human–computer trust scale in a new cultural context.
@@ III-B3 Software Engineering (SE)
Trust is a critical element in SE that influences the adoption of software tools, processes, and collaborative practices. However, our review of 13 articles published between 2002 and 2024 reveals that SE research typically does not engage with established foundational trust models. In fact, only one paper [DBLP:conf/esem/JalaliGS10] cited [Mayer1995]’s trust definition and based their research on a related trust model. Furthermore, one paper [DBLP:journals/ese/WangR16] implicitly referred to established trust models by mentioning risk, that is, the aspect of vulnerability (see Section II-A). The other papers either did not explicitly define trust or used custom definitions, diverging from disciplines such as IS and HCI, where established trust models are more commonly applied.
An example of SE’s customized approach to trust is the PICSE framework proposed by [DBLP:conf/icse/JohnsonBFFZ23] [DBLP:conf/icse/JohnsonBFFZ23]. It describes a structured approach to understanding trust in software tools, identifying five key factors: personal, interaction, control, system, and expectation. The authors refer to [DBLP:conf/chi/FoggT99]’s credibility framework [DBLP:conf/chi/FoggT99] from HCI and [rousseau1998not]’s trust literature review [rousseau1998not] from IS. However, these trust models did not contribute to PICSE but are discussed after introducing the framework.
[DBLP:conf/esem/JalaliGS10] adapted [DBLP:conf/pst/Schultz06]’s situational trust model [DBLP:conf/pst/Schultz06], which extends [Mayer1995]’s model, to analyze trust dynamics in global software engineering [DBLP:conf/esem/JalaliGS10]. In line with HCI and IS research, their model underlines the dynamic nature of trust, distinguishing initial trust building, where trust forms based on past interactions, from trust evolution, where trust fluctuates during the project based on expectations and behaviors.
[DBLP:journals/ese/WangR16] applied evolutionary game theory to examine trust and cooperation in distributed teams. Their study does not explicitly model trust development or draw on established trust theories [DBLP:journals/ese/WangR16]. Similarly, [DBLP:journals/tse/CalinescuWGIHK18] propose dynamic assurance cases for engineering trustworthy self-adaptive software, emphasizing system reliability and correctness rather than interpersonal trust [DBLP:journals/tse/CalinescuWGIHK18]. [DBLP:journals/spe/AkbarKMRD24] take a different approach, presenting a taxonomy of decision-making challenges in trustworthy AI [DBLP:journals/spe/AkbarKMRD24]. Although these studies address trust-related concerns, they do so in domain-specific ways, emphasizing technical, ethical, or structural factors rather than explicitly modeling trust development based on established theories.
[DBLP:journals/ese/BratthallJ02] address trust in single data source software engineering case studies [DBLP:journals/ese/BratthallJ02]. [DBLP:conf/esem/BertramH0WS20] investigate trust perceptions in code review through an eye-tracking study, demonstrating how the perceived origin of code (human versus machine-generated) influences review behaviors and trust judgments [DBLP:conf/esem/BertramH0WS20]. However, both papers do not explicitly reference or build upon established trust models.
@@ III-B4 Literature Reviews and Special Issues
As part of our review, we found five literature reviews of trust that were not limited to just one of our target disciplines SE, HCI, or IS. [DBLP:journals/csur/KaurURD23]’s review centers around requirements for trustworthy AI [DBLP:journals/csur/KaurURD23]. They refer to trust definitions across disciplines, including psychology, sociology, and economy, mentioning integrity and reliability as an agreement across these disciplines. They define trustworthy AI as a framework to ensure that a system is worthy of being trusted based on the evidence concerning its stated requirements [DBLP:journals/csur/KaurURD23]. [DBLP:journals/electronicmarkets/YangW22] examined 142 studies published between 2015 and 2022, identifying components, influencing factors, and outcomes of users’ trust in AI [DBLP:journals/electronicmarkets/YangW22]. [DBLP:journals/csur/LiQLDLPYZ23]’s review outlines aspects of AI trustworthiness such as robustness, explainability, transparency, and fairness, organized along the lifecycle of AI systems [DBLP:journals/csur/LiQLDLPYZ23]. However, they do not discuss established trust definitions or models. [GULATI2024100495] reviewed 47 studies on trust in technology published in IS and HCI venues and found 17 different theories that were integrated from disciplines such as psychology and economics [GULATI2024100495]. They conclude that the intricacies of how trust is formed and maintained in online environments still necessitates further investigation and that the development of standardized and empirically validated instruments for trust measurement is crucial. Finally, [DBLP:conf/hicss/LockeyGHS21]’s review focuses on antecedents of trust in AI [DBLP:conf/hicss/LockeyGHS21]. They identify five AI-specific trust challenges: transparency and explainability, accuracy and reliability, automation versus augmentation, anthropomorphism and embodiment, and mass data extraction [DBLP:conf/hicss/LockeyGHS21]. An interesting observation regarding AI assistants for software development is that over-anthropomorphism may lead to overestimation of the AI’s capabilities [DBLP:conf/hicss/LockeyGHS21], increasing risk [DBLP:journals/chb/CulleyM13], decreasing trust [DBLP:journals/tiis/BakerPUK18], and potentially leading to manipulation [Salles02042020]. This is certainly an underexplored perspective on AI assistants for software development, especially when focusing on human factors of software security.
@@
Three literature reviews focused on trust research in HCI. [DBLP:journals/ijhci/Jeon24]’s review discussed the effects of emotions on trust in the context of technology use [DBLP:journals/ijhci/Jeon24]. Their review indicates that positive emotions lead to higher trust. [DBLP:journals/ijhci/BachKHBS24] reviewed 23 empirical studies that focus on user trust in AI-enabled systems [DBLP:journals/ijhci/BachKHBS24]. They describe trust definitions, factors that influence user trust, and measurement methods, reporting that seven studies explicitly defined trust, six of which referring to [Mayer1995] [Mayer1995] or [Lee2004Trust] [Lee2004Trust]. Eight studies conceptualized trust, while nine studies neither defined nor conceptualized trust. They found surveys, interviews, and focus groups to be the most commonly used methods to assess user trust. [DBLP:journals/ijhci/BachKHBS24] conclude that it is important to select a definition of user trust that aligns with the specific study context. [10.1145/3696449] reviewed 65 articles to assess the various definitions of appropriate trust in human-AI interaction, providing an overview of concepts and definitions and identifying similarities and differences between them [10.1145/3696449]. They conclude that four common methods for building appropriate trust are improving system transparency, considering user cognition and perception, using guidelines or models to achieve calibrated trust in AI, and considering the entire continuum of trust, including over-, under-, mis-, and dis-trust.
@@
One literature review had a clear IS focus. [doi:10.5465/annals.2018.0057] synthesized two decades of empirical research on the determinants of human trust in AI across disciplines [doi:10.5465/annals.2018.0057]. The authors motivate how trust in AI differs from other technologies and identify an AI’s capabilities as an important antecedent of trust development. In addition, they propose a framework that identifies factors that influence users’ cognitive and emotional trust. Regarding the development of cognitive trust, they highlight the importance of an AI’s tangibility, transparency, reliability, and immediacy. For emotional trust, on the other hand, and AI’s anthropomorphism is a major factor. [doi:10.5465/annals.2018.0057] also critically reflect on past studies by pointing to the diversity of trust measure used and the lack of longer-term studies in higher stakes environments [doi:10.5465/annals.2018.0057].
@@
For IS, we further identified three special issues on trust. One of them was published in 2010 in their flagship journal MIS Quarterly, focusing on novel perspectives on trust in IS. The fact that this special issue was published 15 years ago is another indication of the maturity of trust research in that discipline. The other IS-specific special issues focused on trust in online environments [DBLP:journals/jmis/BenbasatGP08] and trust in AI [DBLP:journals/electronicmarkets/LukyanenkoMS22].
@@
For SE, we included two literature reviews. [DBLP:journals/csur/LiuLZJWS25] conducted a tertiary study reviewing 141 secondary studies on trustworthy AI and software, describing trustworthiness as a highly abstract concept comprising related quality attributes [DBLP:journals/csur/LiuLZJWS25]. The review focuses on these quality attributes and compares them between trustworthy AI and other trustworthy software. [DBLP:journals/csur/LiuLZJWS25] do not focus on the conceptualization of trust and trustworthiness. [DBLP:journals/ese/HouJ23] reviewed 112 articles on trust in software ecosystems, identifying different definitions of concepts such as software trust, system trust, or software service trust. However, their focus was on compiling trust definitions, not unifying or aligning them [DBLP:journals/ese/HouJ23].
{results}
RQ2 (trust conceptualizations):
[…]
@@ III-C Critical Review of Trust Conceptualizations in Software Engineering Research
Of the 58 articles that we analyzed as part of our critical review, only one article referred to an established trust model, and only three articles provided trust definitions. This underlines the problematic situation suggested by our literature across disciplines (see Section III-B3). There is a significant gap in the maturity of trust research in SE compared to IS and HCI. Trust as a concept is often discussed in a context-specific manner without embedding concepts and definitions in established trust theories or models. This lack of embedding in established trust theories from other disciplines hinders interdisciplinary comparisons and research toward a deeper understanding of what constitutes trust in SE-specific settings.
@@
[DBLP:conf/icse/NollerS0R22] are the only authors who explicitly reference an established trust model, incorporating [Lee2004Trust]’s model of human-automation trust [DBLP:conf/icse/NollerS0R22]. [DBLP:journals/tse/AlamiPCW22] define trust as the unyielding belief that the person is truthful and reliable and in the context of their research as the willingness of the community to rely on the contributor, a prerequisite for considering her code change [DBLP:journals/tse/AlamiPCW22]. This conceptualization frames trust as a social and reputation-based construct, essential in collaborative environments such as open-source software development, where trust is built through ongoing contributions. Although this framing is reasonable, an embedding in existing trust models would allow researchers to identify the particularities of trust in open-source software compared to other settings.
@@
[DBLP:journals/tse/WinterBCHHNW23] [DBLP:journals/tse/WinterBCHHNW23] implicitly define trust by referencing dispositional trust as proposed by [DBLP:journals/hf/HoffB15] [DBLP:journals/hf/HoffB15], which describes an individual’s inherent tendency to trust automation, independent of context or system performance. This positions trust as a psychological trait that varies between individuals. However, it is important to note that trust was only a peripheral topic, not a central focus of their study. [DBLP:journals/pacmse/KouCW0024] discuss trust in the context of large language models (LLMs) for code generation, suggesting that users assess trustworthiness based on the model’s explainability and how well its predictions align with their expectations [DBLP:journals/pacmse/KouCW0024]. However, they neither explicitly define trust nor consider the broader explainability discussion (see Section II). Instead, they cite previous work by [DBLP:conf/chi/BoggustHSS22], who explored how users determine whether to trust a machine learning model based on its ability to convey behavior in a way that aligns with their expectations [DBLP:conf/chi/BoggustHSS22]. Neither study provides a formal definition of trust; rather, both frame it in terms of user perceptions and interpretability. Given the amount of related work on trustworthiness and explainability of AI-based systems in philosophy (see Section II-C) and IS (see Section III-B1), it is surprising that this article is again not broadly embedded in established trust conceptualizations.
{results}
RQ2 (trust in SE research):
Our critical review confirmed that there is a significant maturity gap in trust-related research in SE compared to IS and HCI. Of the 58 analyzed articles, only one referred to an established trust model, and only three provided trust definitions. This lack of embedding in established trust models hinders interdisciplinary comparisons and research toward a deeper understanding of what constitutes trust in SE-specific settings.